I am trying to figure out the formula I need to use for the following math word problem....
Find the length and width of a rectangle with the length being 3 cm less than twice the width. The parameter is 36.
I know P = 2L+2W
But I can not remember the formula or what I need to add into the formula to get the width then the length.
Thanks

Answer by nerdybill(7384)

Find the length and width of a rectangle with the length being 3 cm less than twice the width. The parameter is 36.
.
Since they don't tell you the width, create a variable to hold it:
Let W = width
then from "the length being 3 cm less than twice the width" we get
2W-3 = length
.
P = 2L+2W
36 = 2(2W-3)+2W
dividing both sides by 2:
18 = (2W-3)+W
18 = 3W-3
21 = 3W
7 cm = W (width)
.
Length:
2W-3 = 2(7)-3 = 14-3 = 11 cm